# Methods and Open Problems in Differentiable Social Choice: Learning Mechanisms, Decisions, and Alignment

### v1

@@ Abstract
Beyond synthesizing methods, this Review is explicitly forward-looking. Across six domains, we identify 36 concrete open problems spanning incentive guarantees, robustness to distribution shift and strategic manipulation, certification and auditability of learned mechanisms, pluralistic preference aggregation, governance of alignment objectives, and the limits of inverse inference from behavior. These open problems delineate a new research agenda at the intersection of machine learning, economics, and democratic theory.
@@ 1 Introduction
Modern deployments have amplified these limitations. Spectrum auctions, online advertising markets, participatory budgeting, decentralized governance platforms, and preference-based training of large language models (LLMs) all operate in decision spaces that are high-dimensional, combinatorial, non-convex, and distribution-dependent. Yet, social choice is (both implicitly and explicitly) used in many of the most important contemporary machine learning research, e.g. in automated auction mechanism design for revenue optimzation Duan et al. (2023) and Large Language Models Duetting et al. (2024), in stake-holder portfolio balance and participatary resource allocation Fairstein et al. (2024), in federated learning where model updates are aggregated based on a certain mechanism Kesari & Kesari (2025), and, most celebrated recently, in preference-based alignment (RLHF) of LLMs from human annotators who have inherently heterogeneous preference models Siththaranjan et al. (2024); Swamy et al. (2024). We summarize these applications of social choice in machine learning in Table 1. In such settings, analytic derivations with universal guarantees can be simultaneously too pessimistic—ruling out mechanisms that perform well on realistic data—and too optimistic—failing catastrophically under distribution shift or strategic adaptation Hart et al. (2013); Caiata et al. (2025). Nevertheless, many contemporary applied systems are already learned: allocation policies, ranking rules, and reward models are trained by gradient descent, whether or not designers acknowledge their implicit social-choice content.
@@
Our goal is not merely to catalog methods, but to articulate a coherent narrative: the transition from axiomatic artifacts to learned institutions. As collective decision-making becomes embedded in machine learning pipelines, the design of loss functions, architectures, and evaluation protocols becomes inseparable from democratic theory. Differentiable social choice thus emerges not only as a technical research program, but as a foundational framework for the governance, auditability, and legitimacy of algorithmic decision-making systems.
@@ 2.2 Differentiable incentive constraints: regret, cyclic monotonicity, and certification
and optimizes expected revenue subject to low expected regret, often via Lagrangian or augmented Lagrangian schemes Dütting et al. (2017). This family of methods is frequently associated with RegretNet-style pipelines, in which an inner maximization (adversarial misreport search) is approximated by gradient ascent on $\hat{v}_{i}$, while an outer minimization updates mechanism parameters Dütting et al. (2017).
@@ 2.6 From auctions to edge markets: partial fulfillment and resource-constrained mechanism design
“Neural Auction Mechanisms for Partial Fulfillment in Edge Computing” proposes architectures explicitly tailored to this domain, decoupling allocation decisions into probability and quantity components (PDNet) and using exchangeable attention mechanisms (EANet) to preserve bidder anonymity while scaling across variable numbers of agents Zhang et al. (2026). The learned mechanisms are trained with revenue objectives and regret-based incentive penalties, while embedding feasibility constraints relevant to edge capacity and service provisioning. A notable contribution is the modularity of the differentiable approach: by altering output parameterizations and constraint projections, the same learning framework can address allocation regimes that are awkward for traditional all-or-nothing auction templates Zhang et al. (2026).
@@ 3.4 Axiomatic auditing and the “mimicry–reasoning” gap
Several mechanisms plausibly contribute to the mimicry–reasoning gap. Many voting rules are piecewise-constant with complex decision boundaries; small perturbations near boundaries can flip winners, making generalization fragile in high-dimensional profile spaces. Moreover, axioms often constrain behavior on rare configurations (e.g., specific paradoxical profiles), so finite-sample mimicry may never “see” the regions where axioms bite. Finally, axioms are global and compositional (they quantify over counterfactual profiles), while supervised training is pointwise. The audit-first perspective thus becomes essential: learned democracies cannot be assumed to inherit normative properties from their labels Hornischer & Terzopoulou (2024).
@@
A complementary response is to explain learned rules in human-legible terms. “Learning to Explain Voting Rules” learns interpretable surrogates (e.g., trees) that approximate voting behavior and can expose which profile features drive outcomes, supporting governance and debugging Kang et al. (2023). This connects neural social choice to the broader ML agenda of post-hoc interpretability, but with unusually crisp normative stakes: explanations are not only for trust but for verifying that the rule embodies the intended democratic compromise.
@@ 3.5 What voting rules actually do in practice: distribution-specific social choice
Implications: personalization, legitimacy, and “bespoke” democracy.
The distribution-specific view suggests a provocative possibility: instead of universal rules, we may design electorate-adapted rules that are “more democratic” for a given society (in the sense of fewer violations of the society’s chosen axioms under its empirical preference structure). This resonates with empirical social choice and preference restriction models (e.g., single-peaked, single-crossing, Mallows) that already show how structure changes both feasibility and guarantees Marden (1995); Elkind et al. (2016). However, it also raises legitimacy questions: if rules are trained on historical data, which populations are represented, how robust is the rule to demographic change, and who gets to specify the axiom weights?
@@ 3.7 Differential Voting: loss functions as voting rules
Differential Voting proposes smooth surrogates for discontinuous criteria such as Copeland (pairwise majority wins) and Kemeny (minimizing pairwise disagreement), and establishes consistency and limiting behavior as smoothing vanishes An et al. (2026). Conceptually, this turns “choose a voting rule” into “choose a loss” and enables mix-and-match democracy: weighted sums of losses can interpolate between ideals (e.g., Condorcet consistency vs. smooth optimization stability). This approach also clarifies a central methodological tension in neural mechanism design: objectives that are normatively attractive often induce non-smooth or combinatorial optimization, while smooth objectives may bake in unintended normative biases.
@@ 4 Participatory Budgeting and Complex Allocation
Table 3 summarizes classical and learned participatory budgeting pipelines, highlighting which properties (representation, feasibility, and auditability) are guaranteed by design versus only evaluated empirically.
@@ 4.6 Beyond PB: complex allocation as differentiable social choice
The second is particularly subtle: a model can optimise a fairness proxy (e.g., satisfaction Gini) while systematically failing coalition-based representation axioms that matter for democratic legitimacy Peters et al. (2021). This motivates “axiomatic evaluation suites” for PB analogous to those developed for neural voting rules, but adapted to cost constraints and portfolio outputs.
@@ 5.2 Strategic incentives and endogenous delegation
The tension between epistemic benefits (delegating to experts) and democratic concerns (power concentration) has motivated formal influence measures and constraints. “Rethinking Influence in Liquid Democracy” argues that static delegation analysis misses the dynamic nature of influence accumulation and studies how influence should be measured and controlled over time Grossi & Nitsche (2025). This perspective aligns closely with FL, where influence concentration can correspond to over-weighting a small set of clients, risking bias and poisoning vulnerability.
@@ 5.5 From liquid voting to federated learning: fluid delegation as decentralized aggregation
Viscous-retained democracy (FedVRD): delegation for bandwidth and robustness.
“Fluid Democracy in Federated Data Aggregation” explicitly maps LD topology onto FL: clients evaluate neighbor competence (e.g., via local validation loss) and delegate their model updates to more competent peers in a peer-to-peer graph, reducing direct communication with the server Kesari & Kesari (2025). The proposed protocol, viscous-retained democracy (FedVRD), introduces a “viscosity” parameter that dampens influence flow, aiming to preserve epistemic gains (routing updates through competent nodes) without allowing unbounded influence accumulation Kesari & Kesari (2025). The paper emphasizes two practical payoffs: (i) communication savings by reducing the number of server-bound uploads, and (ii) improved robustness to poisoning, as viscous delegation can dilute adversarial influence and limit amplification through delegation chains Kesari & Kesari (2025).
Relation to robust and secure aggregation.
FedVRD is complementary to robust aggregation rules developed for Byzantine-resilient FL. Classical defenses include Krum and its variants, which select updates close to the majority in Euclidean space Blanchard et al. (2017), trimmed mean/coordinate-wise median aggregators, and cryptographic secure aggregation protocols that hide individual updates from the server Bonawitz et al. (2017). The LD-inspired shift is architectural: rather than relying purely on server-side filtering, delegation places part of the aggregation burden on the network, enabling hierarchical or multi-hop aggregation. This resembles communication-efficient FL designs (client selection, hierarchical FL) but adds a trust graph as a principled organizing primitive Kesari & Kesari (2025).
@@ 6.1 Section overview: the hidden ballot in the loss function
Modern alignment pipelines for large language models (LLMs) are, at their core, preference aggregation systems. Reinforcement Learning from Human Feedback (RLHF) learns from many pairwise comparisons—“output $A$ is better than output BB”—collected from annotators who may disagree for principled, contextual, or idiosyncratic reasons Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022). Direct Preference Optimization (DPO) and related “direct alignment” objectives bypass explicit reward modeling but still optimize a loss induced by preference data Rafailov et al. (2023). In both cases, the training procedure compresses heterogeneous human judgments into a single reward model or policy, thereby performing an implicit social choice operation.
@@
Majority
resp.
Pareto
IIA/LIIA
Condorcet
(A) Standard alignment objectives as implicit voting rules
(C) Axiomatic loss design (Differential Voting)
@@ 8.1 From axiomatic artifacts to learned institutions
This Review has traced a coherent paradigm shift: from axiomatic democracy, where mechanisms are derived as static artifacts of logic, to differentiable social choice, where mechanisms are implemented as dynamic, parameterized functions optimized from data Dütting et al. (2017; 2019); Anil & Bao (2021); Fairstein et al. (2024); Kesari & Kesari (2025); An & Du (2026). The practical motivation is clear. Classical social choice and mechanism design have produced profound conceptual tools and impossibility results—Arrow, Gibbard–Satterthwaite, and Myerson remain defining pillars Arrow (2020); Gibbard (1973); Satterthwaite (1975); Myerson (1981). Yet in many modern deployments (multi-item auctions, constrained public spending, decentralized edge aggregation, and preference-based training of frontier models), the relevant decision spaces are high-dimensional, non-convex, and distribution-dependent. In these regimes, the classical program of closed-form derivations and universal guarantees can be simultaneously too rigid (unable to exploit structure in real data) and too pessimistic (worst-case failures that may be vanishingly rare or, conversely, dangerously frequent) Hart et al. (2013); Caiata et al. (2025).
@@ 8.2 Cross-cutting methodological lessons
A recurring theme is that the most important failures often arise under misspecification: shifted bidder populations, new preference cultures, adversarial manipulation, and changing platform dynamics. The economics literature has begun to formalize these concerns via refined notions of distributional robustness that avoid “performance cliffs” when the ambiguity set is slightly wrong Ball & Kattwinkel (2024). In social choice, the analogous move is from binary axiom satisfaction to empirical violation rates under realistic distributions, which can invert classical intuitions about which rule is “more democratic” in practice Caiata et al. (2025). In security-focused work, robustness becomes adversarial training against manipulative coalitions or strategic learners Li et al. (2026); Nedelec et al. (2019).

### v2

@@ Abstract
Beyond synthesizing methods, this Review is explicitly forward-looking. Across six domains, we identify 36 concrete open problems spanning incentive guarantees, robustness to distribution shift and strategic manipulation, certification and auditability of learned mechanisms, pluralistic preference aggregation, governance of alignment objectives, and the limits of inverse inference from behavior. These open problems delineate a new research agenda at the intersection of machine learning, economics, and social choice theory.
@@ 1 Introduction
Modern deployments have amplified these limitations. Spectrum auctions, online advertising markets, participatory budgeting, decentralized governance platforms, and preference-based training of large language models (LLMs) all operate in decision spaces that are high-dimensional, combinatorial, non-convex, and distribution-dependent. Yet, social choice is (both implicitly and explicitly) used in many of the most important contemporary machine learning research, e.g. in automated auction mechanism design for revenue optimzation Duan et al. (2023) and Large Language Models joint creative generations Duetting et al. (2024), in stake-holder portfolio balance and participatary resource allocation Fairstein et al. (2024), in federated learning where model updates are aggregated based on a certain mechanism Kesari & Kesari (2025), and, most celebrated recently, in preference-based alignment (RLHF) of LLMs from human annotators who have inherently heterogeneous preference models Siththaranjan et al. (2024); Swamy et al. (2024). We summarize these applications of social choice in machine learning in Table 1. In such settings, analytic derivations with universal guarantees can be simultaneously too pessimistic—ruling out mechanisms that perform well on realistic data—and too optimistic—failing catastrophically under distribution shift or strategic adaptation Hart et al. (2013); Caiata et al. (2025). Nevertheless, many contemporary applied systems are already learned: allocation policies, ranking rules, and reward models are trained by gradient descent, whether or not designers acknowledge their implicit social-choice content.
@@
Our goal is not merely to catalog methods, but to articulate a coherent narrative: the transition from axiomatic artifacts to learned institutions. As collective decision-making becomes embedded in machine learning pipelines, the design of loss functions, architectures, and evaluation protocols becomes inseparable from social choice theory. Differentiable social choice thus emerges not only as a technical research program, but as a foundational framework for the governance, auditability, and legitimacy of algorithmic decision-making systems.
@@ 2.2 Differentiable incentive constraints: regret, cyclic monotonicity, and certification
where $u_{i}(\hat{v}_{i},v_{-i};\theta)$ denotes agent $i$’s quasi-linear utility under a misreport $\hat{v}_{i}$ when the other agents report $v_{-i}$ and the mechanism is parameterized by $\theta$, while $u_{i}(v_{i},v_{-i};\theta)$ is the utility under truthful reporting; the regret $\mathrm{Regret}_{i}(\theta;v)$ therefore measures the maximum utility gain achievable by unilateral deviation from truthfulness at valuation profile $v$. Regret-based training optimizes expected revenue subject to low expected regret, often via Lagrangian or augmented Lagrangian schemes Dütting et al. (2017). This family of methods is frequently associated with RegretNet-style pipelines, in which an inner maximization (adversarial misreport search) is approximated by gradient ascent on $\hat{v}_{i}$, while an outer minimization updates mechanism parameters Dütting et al. (2017).
@@ 2.6 From auctions to edge markets: partial fulfillment and resource-constrained mechanism design
Neural Auction Mechanisms for Partial Fulfillment in Edge Computing proposes architectures explicitly tailored to this domain, decoupling allocation decisions into probability and quantity components (PDNet) and using exchangeable attention mechanisms (EANet) to preserve bidder anonymity while scaling across variable numbers of agents Zhang et al. (2026). The learned mechanisms are trained with revenue objectives and regret-based incentive penalties, while embedding feasibility constraints relevant to edge capacity and service provisioning. A notable contribution is the modularity of the differentiable approach: by altering output parameterizations and constraint projections, the same learning framework can address allocation regimes that are awkward for traditional all-or-nothing auction templates Zhang et al. (2026).
Figure 2: Comparison of Classical Ad Auction vs. LLM Joint Creative Mechanisms. On the left, a classical ad auction ranks advertisers by bid times quality to allocate fixed slots. On the right, LLM agents representing advertisers produce next-token distributions that are aggregated via a differentiable token auction mechanism. The aggregation proceeds token-by-token with bids influencing the combined output distribution; the resulting creative jointly reflects advertiser preferences and supports gradient-based learning.
Complementing these neural mechanisms in edge settings, Mechanism Design for Large Language Models Duetting et al. (2024) proposes a token auction model for aggregating outputs from multiple self-interested LLM agents in joint content generation. Rather than allocating discrete items, this framework operates on a token-by-token basis, letting each LLM agent submit a single-dimensional bid that influences the aggregated next-token distribution in generation. We illustrate the process and compare it with classical Ad Auction process in Figure 2. Duetting et al. (2024) formalize incentive properties for such mechanisms and show they are equivalent to a monotonicity condition on output aggregation, enabling analogues of second-price payments even without explicit valuation functions Duetting et al. (2024). They further derive concrete aggregation rules — e.g., bid-weighted averages in linear and log-space — that align joint generation outcomes with agents’ implicit preferences over distributions. This token auction approach not only bridges classical auction design with generative AI but also provides a differentiable mechanism that can be integrated with learning-based policies, illustrating how economic mechanisms and gradient-based optimization can jointly shape multi-agent LLM outputs.
@@ 3.4 Axiomatic auditing and the “mimicry–reasoning” gap
Several mechanisms plausibly contribute to the mimicry–reasoning gap. Many voting rules are piecewise-constant with complex decision boundaries; small perturbations near boundaries can flip winners, making generalization fragile in high-dimensional profile spaces. Moreover, axioms often constrain behavior on rare configurations (e.g., specific paradoxical profiles), so finite-sample mimicry may never “see” the regions where axioms bite. Finally, axioms are global and compositional (they quantify over counterfactual profiles), while supervised training is pointwise. The audit-first perspective thus becomes essential: learned social choice cannot be assumed to inherit normative properties from their labels Hornischer & Terzopoulou (2024).
@@
A complementary response is to explain learned rules in human-legible terms. “Learning to Explain Voting Rules” learns interpretable surrogates (e.g., trees) that approximate voting behavior and can expose which profile features drive outcomes, supporting governance and debugging Kang et al. (2023). This connects neural social choice to the broader ML agenda of post-hoc interpretability, but with unusually crisp normative stakes: explanations are not only for trust but for verifying that the rule embodies the intended compromise.
@@ 3.7 Differential Voting: loss functions as voting rules
Differential Voting proposes smooth surrogates for discontinuous criteria such as Copeland (pairwise majority wins) and Kemeny (minimizing pairwise disagreement), and establishes consistency and limiting behavior as smoothing vanishes An et al. (2026). Conceptually, this turns “choose a voting rule” into “choose a loss” and enables mix-and-match social choice: weighted sums of losses can interpolate between ideals (e.g., Condorcet consistency vs. smooth optimization stability). This approach also clarifies a central methodological tension in neural mechanism design: objectives that are normatively attractive often induce non-smooth or combinatorial optimization, while smooth objectives may bake in unintended normative biases.
@@ 4 Participatory Budgeting and Complex Allocation
Table 3 summarizes classical and learned participatory budgeting pipelines, highlighting which properties (representation, feasibility, and auditability) are guaranteed by design versus only evaluated empirically. Figure 3 illustrates the transition from classical, rule-based participatory budgeting to LLM-enabled and differentiable approaches that learn portfolio allocations under budget and normative constraints.
Figure 3: From axiomatic participatory budgeting to learned aggregation over portfolios. The left panel illustrates classical participatory budgeting, where citizens submit ballots over public projects with fixed costs, and a hand-designed rule (e.g., greedy knapsack, Phragmén, or the Method of Equal Shares) selects a feasible project portfolio under a hard budget constraint, inheriting specific axiomatic guarantees such as proportionality and representation. The right panel depicts LLM-based and differentiable participatory budgeting, where richer preference inputs—including natural-language deliberation—are mapped to feasible portfolios by a learned aggregation model trained end-to-end against target normative metrics (e.g., fairness, minority coverage, efficiency, interpretability), reframing rule choice as continuous optimization over social objectives rather than ex-ante commitment to a single aggregation rule.
@@ 4.6 Beyond PB: complex allocation as differentiable social choice
The second is particularly subtle: a model can optimise a fairness proxy (e.g., satisfaction Gini) while systematically failing coalition-based representation axioms Peters et al. (2021). This motivates “axiomatic evaluation suites” for PB analogous to those developed for neural voting rules, but adapted to cost constraints and portfolio outputs.
@@ 5 Liquid Democracy and Decentralized Aggregation
Figure 4 illustrates how liquid democracy provides a unifying abstraction for delegation and aggregation, bridging classical social choice with modern graph-based aggregation mechanisms used in federated learning.
Figure 4: Liquid democracy as a bridge between social choice and federated learning. Left: In classical liquid democracy, voters either cast ballots directly or delegate their voting power along a directed delegation graph, inducing transitive power flow, concentration, and potential cycles. Right: The same delegation structure can be repurposed in federated learning, where clients compute local updates and influence propagates through a trust or delegation graph to determine weighted global aggregation. Across both settings, the delegation graph is not merely an input but a design variable, enabling differentiable, adaptive, and robust aggregation mechanisms that connect social choice theory with learning-based systems.
@@ 5.2 Strategic incentives and endogenous delegation
The tension between epistemic benefits (delegating to experts) and axiomatic concerns (power concentration) has motivated formal influence measures and constraints. “Rethinking Influence in Liquid Democracy” argues that static delegation analysis misses the dynamic nature of influence accumulation and studies how influence should be measured and controlled over time Grossi & Nitsche (2025). This perspective aligns closely with FL, where influence concentration can correspond to over-weighting a small set of clients, risking bias and poisoning vulnerability.
@@ 6 AI Alignment as a Social Choice Problem
Figure 5: AI alignment as a social choice problem. Human annotators provide heterogeneous and often conflicting pairwise judgments over candidate model outputs (left), which are aggregated by alignment objectives such as RLHF or DPO through a preference-based loss function that acts as a hidden ballot box (center), implicitly corresponding to a particular social choice rule (e.g., BTL-induced Borda aggregation); optimization then collapses this plural input into a single trained policy and output (right), importing classical social-choice pathologies—such as majority insensitivity, context erasure, and vulnerability to manipulation—into alignment pipelines.
@@ 6.1 Section overview: the hidden ballot in the loss function
Modern alignment pipelines for large language models (LLMs) are, at their core, preference aggregation systems. Reinforcement Learning from Human Feedback (RLHF) learns from many pairwise comparisons—“output $A$ is better than output BB”—collected from annotators who may disagree for principled, contextual, or idiosyncratic reasons Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022). Direct Preference Optimization (DPO) and related “direct alignment” objectives bypass explicit reward modeling but still optimize a loss induced by preference data Rafailov et al. (2023). In both cases, the training procedure compresses heterogeneous human judgments into a single reward model or policy, thereby performing an implicit social choice operation. Figure 5 illustrates how modern alignment objectives implement an implicit social choice rule, compressing heterogeneous human preferences into a single optimized behavior via the loss function rather than discovering a uniquely “correct” reward.
@@ 6.6 Strategic feedback, poisoning, and mechanism design for alignment pipelines
Majority
resp.
Pareto
IIA/LIIA
Condorcet
(A) Standard alignment objectives as implicit voting rules
(C) Axiomatic loss design (Differential Voting)
@@ 8.1 From axiomatic artifacts to learned institutions
This Review has traced a coherent paradigm shift: from axiomatic social choice, where mechanisms are derived as static artifacts of logic, to differentiable social choice, where mechanisms are implemented as dynamic, parameterized functions optimized from data Dütting et al. (2017; 2019); Anil & Bao (2021); Fairstein et al. (2024); Kesari & Kesari (2025); An & Du (2026). The practical motivation is clear. Classical social choice and mechanism design have produced profound conceptual tools and impossibility results—Arrow, Gibbard–Satterthwaite, and Myerson remain defining pillars Arrow (2020); Gibbard (1973); Satterthwaite (1975); Myerson (1981). Yet in many modern deployments (multi-item auctions, constrained public spending, decentralized edge aggregation, and preference-based training of frontier models), the relevant decision spaces are high-dimensional, non-convex, and distribution-dependent. In these regimes, the classical program of closed-form derivations and universal guarantees can be simultaneously too rigid (unable to exploit structure in real data) and too pessimistic (worst-case failures that may be vanishingly rare or, conversely, dangerously frequent) Hart et al. (2013); Caiata et al. (2025).
@@ 8.2 Cross-cutting methodological lessons
A recurring theme is that the most important failures often arise under misspecification: shifted bidder populations, new preference cultures, adversarial manipulation, and changing platform dynamics. The economics literature has begun to formalize these concerns via refined notions of distributional robustness that avoid “performance cliffs” when the ambiguity set is slightly wrong Ball & Kattwinkel (2024). In social choice, the analogous move is from binary axiom satisfaction to empirical violation rates under realistic distributions Caiata et al. (2025). In security-focused work, robustness becomes adversarial training against manipulative coalitions or strategic learners Li et al. (2026); Nedelec et al. (2019).